# Towards Building Multilingual Language Model for Medicine

### v2

@@ Abstract
In this paper, we aim to develop an open-source, multilingual language model for medicine, that the benefits a wider, linguistically diverse audience from different regions. In general, we present the contribution from the following aspects: first, for multilingual medical-specific adaptation, we construct a new multilingual medical corpus, that contains approximately 25.5B tokens encompassing 6 main languages, termed as MMedC, that enables auto-regressive training for existing general LLMs. second, to monitor the development of multilingual LLMs in medicine, we propose a new multilingual medical multi-choice question-answering benchmark with rationale, termed as MMedBench; third, we have assessed a number of popular, open-source large language models (LLMs) on our benchmark, along with those further auto-regressive trained on MMedC, as a result, our final model, termed as MMedLM 2, with only 7B parameters, achieves superior performance compared to all other open-source models, even rivaling GPT-4 on MMedBench. We will make the resources publicly available, including code, model weights, and datasets.
@@ 1 Introduction
In the recent literature, large language models (LLMs) have demonstrated great promise in healthcare, for example, closed-source models such as GPT-4 [20] and MedPalm-2 [23] have shown remarkable performance, and successfully passed the United States Medical Licensing Examination (USMLE). Concurrently, open-source models like Llama 2 have also facilitated the development of specialized language models for medicine, such as PMC-LLaMA, MedAlpaca, and ChatDoctors [29, 6, 31], gradually bridging the performance gap with their closed-source peers. Despite these advancements, the primary focus on English-language applications by these sophisticated medical language models has constrained their potential reach, limiting the benefits to a wider, linguistically diverse audience.
Figure 1: Overview of our contributions. Figure a demonstrates our proposed large-scale multilingual medical corpus (MMedC), containing 25.5B tokens, covering six main language, collected from four data sources. Figure b shows the composition our comprehensive multilingual medical benchmark (MMedBench), that is constructed by aggregating medical QA cases in different languages, and prompting GPT-4 to provide rationale sentences. MMedBench enables the evaluation on both multi-choice accuracy and the ability of rationale generation for different LLMs under zero-shot or fine-tuning settings. In Figure c left, the final multi-choice accuracy of various LLMs on our MMedBench are show, where our final model MMedLM 2 demonstrated the best performance among all existing open-source LLMs. In Figure c right, we further detail the gains in both multi-choice accuracy and ability of rationale generation, when comparing MMedLM to InternLM, or comparing MMedLM 2 to InternLM 2. Considering that the main difference between the MMedLM series and the InterLM series lies in the auto-regressive training on MMedC, such comparison highlights the importance of our contributed medical-specific multilingual language corpus.
@@
For auto-regressive training, we have developed a large-scale Multilingual Medical Corpus (MMedC), amassing over 25.5 billion medical-related tokens across six primary languages: English, Chinese, Japanese, French, Russian, and Spanish. This diverse dataset was compiled from four distinct sources: (i), we devised an automatic pipeline to filter medical-related content from the broad multilingual corpus, ensuring a focused and relevant dataset, (ii), we curated an extensive collection of medical textbooks in various languages, and converted them into texts with carefully designed pre-processing, e.g., Optical Character Recognition (OCR), heuristic data filtering, etc. Due to copyright issue, we will share the name list of the books and the methodologies and tools for curation, (iii), to guarantee a wide-ranging encapsulation of medical knowledge, we incorporated texts from some famous open-source medical websites, enriching our corpus with authoritative and comprehensive medical information, (iv), additionally, we have integrated a number of existing small-scale medical corpus datasets, further enhancing the breadth and depth of ours. To our knowledge, MMedC represents the first endeavor to construct a corpus specifically focused on the multilingual medical domain.
@@
As for benchmark curation, we start by aggregating existing medical multiple-choice QA datasets across the six languages as of MMedC. We further augment them with rationale content by using GPT-4, enriching the datasets with explanations that support the correct answers. Consequently, our enriched dataset encompasses 53,566 QA pairs across six languages, uniquely offering both multi-choice QA and accompanying rationale reasoning. This extensive collection spans 21 medical fields, including but not limited to Internal Medicine, Biochemistry, Pharmacology, and Psychiatry, among others, termed as the Multilingual Medical Benchmark (MMedBench). We divide it into 45,048 training pairs and 8,518 testing pairs. The training split enables to finetune LLMs after domain-specific fine-tuning. We utilize the entire test set, comprising 8,518 QA pairs, to evaluate the accuracy of multi-choice question answering. To further examine the models’ reasoning ability, we select a subset of 1,136 QA pairs, each accompanied by manually verified rationale sentences, serving as a more specialized benchmark for reasoning evaluation.
@@
At the evaluation phase, we conducted comprehensive benchmarking across eleven existing LLMs with multilingual support, including, GPT-3.5, GPT-4, Gemini-1.0 pro, BLOOM, InternLM, InternLM 2, MedAlpaca, ChatDoctor, PMC-LLaMA, and Mistral, alongside the LLMs further trained with MMedC. These models were evaluated across three different settings: zero-shot, parameter-efficient fine-tuning (PEFT), and full fine-tuning. Given the complexity of evaluating rationale quality, which demands an assessment of long sentence semantic integrity, in addition to leveraging mainstream automated metrics, we also incorporate human rating scores in our analysis. This dual approach not only provides a comprehensive measure on each model’s performance, but also enables us to scrutinize the correlation between automated metrics and human judgment. Through this analysis, we identify the most reliable metric for extended comparisons, thereby enriching the methodology for evaluating reasoning ability in large language models.
@@
In our experiments, models that underwent further auto-regressive training on the MMedC consistently demonstrate enhanced performance, thereby underscoring the value and effectiveness of our compiled multilingual corpus. We will publicly release our dataset (except for the license-restricted books for which we will provide a names list), codebase, and trained models to foster future research. In addition, we recognize the significance of robust evaluation metrics, particularly for the generation of medical texts that often involve complex, long sentences. To this end, we will also release detailed human rating results for individual cases.
@@ 2 Results
Here, we start by presenting the statistics of our constructed datasets. Then, we evaluate various LLMs on MMedBench on multi-choice questions and rationale ability, as well as verifying the effectiveness of MMedC. Lastly, we conduct a series of ablation studies to investigate deeper on the impact of each data phase.
2.1 Data Statistics
@@
2.1.1 Multilingual Medical Corpus (MMedC)
MMedC refers to a multilingual medical corpus, that contains over 25.5B tokens, acquired mainly from four sources, i.e., filtering medical-related content from the general large-scale multilingual corpus, medical textbooks, medical websites and existing small-scale corpus. The main statistic results are plotted in Figure 2.
@@ 2.1.1 Multilingual Medical Corpus (MMedC)
Figure 2: Statistic results on MMedC. Figure (a) shows our collected corpora can cover most main countries worldwide. Figure (b) shows the detail token number for different languages and Figure (c) shows how the four considered data sources contributes for different languages, i.e., filtering content, medical textbooks, medical websites and small-scale corpus.
2.1.2 Multilingual Medical Benchmark (MMedBench)
@@
MMedBench represents a comprehensive multilingual medical Question and Answering (QA) benchmark. We start by delving into its core attributes, including the total number of training and testing cases, the distribution of answer options, and the average length of question-and-answer tokens. Figure 3a illustrates these fundamental characteristics, highlighting that MMedBench often includes questions with multiple correct options, that brings complexity for models to navigate. Additionally, the answers contain rationale sections averaging 200 tokens each. This substantial token count serves a dual purpose: it not only aids in training language models by exposing them to extended reasoning passages but also in evaluating their ability to generate and understand lengthy, complex reasoning statements.
@@
In our detailed exploration of MMedBench, we categorize each question into one of 21 medical topic categories using GPT-4. These categories include Internal Medicine, Biochemistry, Pharmacology, Psychiatry, Microbiology, Physiology, Pathology, Immunology, Obstetrics and Gynecology, Public Health, Hematology, Surgery, Emergency Medicine, Orthopedics, Neurology, Anatomy, Medical Genetics, Radiology, Dermatology, and Endocrinology. This categorization has been rigorously verified by at least two clinicians to ensure its comprehensiveness, covering the breadth of medical disciplines.
@@
Figure 3b showcases the diversity of our multilingual benchmark, spanning a wide array of medical questions from foundational clinical medicine to specialized areas such as pharmacology and public health, with a pronounced emphasis on areas like Internal Medicine and Biochemistry. This underlines the benchmark’s capacity to assess a model’s ability to recognize and process a broad spectrum of medical inquiries effectively.
@@
Figure 3: Statistic results on MMedBench. Figure a shows the basic statistic number on the train and test set of MMedBench. The term "Avg. tokens" represents the mean token length per sample across various compositions in it. “Rationale” denotes the rationale sentences in answer. “Option” denotes the option descriptions in choice list and “question” denotes the question sentences. Then the term “Prop. of multi-option” denotes the proportion of the question with multiple correct options and “Prop. of single-option” denotes the proportion of those with one options in answer. The final term “Number of QA pairs” denotes how many QA pairs are in train or test split. Figure b shows the topics distribution in the test split of MMedBench, covering a wide range of medical aspects, ranging from general and specialized medicine to basic medical sciences. This allows MedQA to comprehensively measure the performance of medical models.
2.2 Evaluation on MMedBench
@@
In this section, we present a comprehensive benchmark of the foremost LLMs using our MMedBench under zero-shot, PEFT, and full fine-tuning settings. Our evaluation focuses on two aspects of model performance: the accuracy in multiple-choice questions and the models’ ability to generate rationales. The evaluated LLMs can be categorized into four distinct classes, i.e., close-source LLMs, popular open-source LLMs, medical-specific open-source LLMs, and the those further undergone training on our MMedC. The later three can all be categorized into open-source LLMs.
@@
Initially, our analysis focuses on the state-of-the-art, proprietary close-source LLMs developed by OpenAI and Google, specifically GPT-3.5, GPT-4, and Gemini-1.0 pro. These models are examined through their publicly available online API solely in a zero-shot setting, as they are not accessible for any further training. However, note that, as the training data for these close-source models are confidential, it is difficult to judge whether they are really “zero-shot”. Following this, our evaluation encompasses a range of open-source LLMs such as Llama 2, Mistral, and InternLM 2. We observe that the response from these open-source LLMs to instructions is relatively poor, making it difficult to draw effective comparisons (check Supplementary Material D for more zero-shot fail cases) in the zero-shot setting. We therefore compare them in fine-tuning settings (PEFT and full fine-tuning). Among these, we make a further distinction between general LLMs and those specifically tailored for the medical domain. Finally, we evaluate models that have undergone further training on our proposed corpus, named MMedLM (based on InternLM) and MMedLM 2 (based on InternLM 2). These models are uniquely augmented with domain-specific knowledge, achieved by auto-regressive training on the MMedC.
2.2.1 On Multilingual Multiple-choice Question & Answering
@@
As illustrated in Table 1, medical-specific Large Language Models (LLMs) generally exhibit high accuracy scores in English, yet their performance significantly declines in languages other than English. Notably, the finetuned PMC-LLaMA achieved an English accuracy score of 47.53, despite outperforming its contemporaneous counterparts, falls behind the GPT models significantly. Later, with the deployment of more advanced foundational models, open-source models started to bridge the gap with the GPT series, for instance, Mistral and InternLM 2, once fine-tuned on the train set of MMedBench, recorded average accuracy scores of 60.73 and 58.59, respectively, surpassing all predecessors of comparable scale. Enhanced performance is also observed after additional auto-regressive training on our proprietary MMedC dataset. Specifically, both MMedLM and MMedLM 2, built on top of InternLM and InternLM 2 respectively, demonstrated significant improvements over their counterparts without training on MMedC, for example, 67.30 (MMedLM 2) vs. 58.59 (InternLM 2) under full fine-tuning evaluation. A similar story holds for the PEFT setting, i.e., later LLMs performs better and training on MMedC brings a significant gain. As a result, MMedLM 2 refers to the most competitive open-source model with a 7B parameter count in proximity to GPT-4’s accuracy of 74.27.
@@
Parameter-efficient Fine-tuning Evaluation
@@
MMedLM 2 (Ours) 7B - trainset 61.74 80.01 61.81 52.09 80.47 67.65 67.30
2.2.2 Machine Rating on Rationale
@@
We offer detailed instructions that prompt the model to outline its analytical process for delivering the final answer, enabling a clear assessment of its reasoning capabilities. The performance is then meticulously evaluated using a variety of metrics. Specifically, the ROUGE-1 and BLEU-1 scores are presented in Table 2. Additionally, results for other metrics are detailed in Supplementary Material C providing a comprehensive view of the model’s performance across diverse evaluation frameworks.
@@
MMedLM 41.63/ 38.83 44.30/ 46.38 38.61/ 46.90 37.54/ 37.78 19.99/ 21.28 40.79/ 38.77 37.14/ 38.32
MMedLM 2 47.07/ 41.51 47.15/ 48.36 47.90/ 52.24 43.22/ 41.36 27.81/ 25.70 46.17/ 42.64 43.22/ 41.97
@@
2.2.3 Human Rating on Rationale.
@@
Specifically, from the test set of MMedBench, we randomly selected 50 test cases per language, alongside outcomes generated by six notable models: our MMedLM, InternLM, BLOOM, LLaMA 2, PMC-LLaMA, and GPT-3.5. The sequence of samples and corresponding model outputs was randomized to prevent bias. The evaluation panel, comprising five post-graduate students from the medical school of Shanghai Jiao Tong University and Peking Union Medical College, was instructed to rank the outputs based on accuracy, reasoning ability, and internal knowledge. To facilitate accurate assessment, manually verified references were also provided. Rankings were quantitatively assigned, with the highest rank awarded a score of 6 and the lowest a score of 1, thereby quantifying the quality of each model’s output. In parallel, we leveraged GPT-4 as an additional evaluator, assigning it the role of a judge to rank the outputs. Further details on GPT-4’s evaluation method are available in Supplementary Material A.
@@
Figure 4a illustrates the comparative analysis of model performances through relative ratings. Notably, MMedLM achieved the highest scores in both human (4.54) and GPT-4 (4.39) evaluations, aligning with its superior performance as indicated by the automatic machine metrics. Interestingly, GPT-3.5 received a lower human rating of 3.08, suggesting that the evaluators’ preferences might be influenced by the brevity of the responses. Comprehensive rating results for each language and model are detailed in Supplementary Material C. Given that the rating metric is a relative indicator, adding new models requires a complete re-evaluation process, which is time-consuming due to extensive human labor. Consequently, the latest LLMs such as Mistral, InternLM 2, and our MMedLM 2 have not participated in rating comparisons.
@@
In addition to comparing different LLMs, our study delves into the correlation between various automatic evaluation metrics and human preferences. This correlation analysis enables us to identify the most effective automatic metric for benchmarking purposes, thereby potentially eliminating the need for resource-intensive human evaluations in future research. We employ the Kendall rank correlation coefficient to measure the agreement between the rankings of each model’s generated rationales by automatic metrics and human evaluations. The findings, illustrated in Figure 4b, indicate that GPT-4’s evaluation results have the highest correlation with human judgments, with a $\tau$ value of 0.555. However, it’s important to note that GPT-4’s ratings, while highly correlated, are relative and not easy to scale for evaluating newly introduced models. Among the absolute automatic metrics, ROUGE-1 emerged as the most reliable indicator, demonstrating a $\tau$ value of 0.512. For BLEU-based metrics, BLEU-1 proved to be the most predictive. Consequently, we advocate for the use of ROUGE-1 and BLEU-1 as benchmarks for assessing the rationale capabilities of newly introduced LLMs on MMedBench in subsequent studies.

### v3

@@ Abstract
The development of open-source, multilingual medical language models can benefit a wide, linguistically diverse audience from different regions. To promote this domain, we present contributions from the following: First, we construct a multilingual medical corpus, containing approximately 25.5B tokens encompassing 6 main languages, termed as MMedC, enabling auto-regressive domain adaptation for general LLMs; Second, to monitor the development of multilingual medical LLMs, we propose a multilingual medical multi-choice question-answering benchmark with rationale, termed as MMedBench; Third, we have assessed a number of open-source large language models (LLMs) on our benchmark, along with those further auto-regressive trained on MMedC. Our final model, MMed-Llama 3, with only 8B parameters, achieves superior performance compared to all other open-source models on both MMedBench and English benchmarks, even rivaling GPT-4. In conclusion, in this work, We present a large-scale corpus, a benchmark and a series of models to support the development of multilingual medical LLMs.
@@ 1 Introduction
In the recent literature, large language models (LLMs) have demonstrated great promise in healthcare, for example, closed-source models such as GPT-4 [1] and MedPalm-2 [38] have shown remarkable performance, and successfully passed the United States Medical Licensing Examination (USMLE). Concurrently, open-source models like Llama 2 have also facilitated the development of specialized language models for medicine, such as MEDITRON, PMC-LLaMA, MedAlpaca, and ChatDoctors [48, 11, 25, 6], gradually bridging the performance gap with their closed-source peers. Despite these advancements, the primary focus on English-language applications by these sophisticated medical language models has constrained their potential reach, limiting the benefits to a wider, linguistically diverse audience.
Figure 1: Overview of our contributions. a The figure demonstrates our proposed large-scale multilingual medical corpus (MMedC), containing 25.5B tokens, covering six main languages, collected from four data sources. b The figure shows the composition of our comprehensive multilingual medical benchmark (MMedBench), that is constructed by aggregating medical QA cases in different languages, and prompting GPT-4 to provide rationale sentences. MMedBench enables the evaluation on both multi-choice accuracy and the ability of rationale generation for different LLMs under zero-shot or fine-tuning settings. c The line plot shows the final multi-choice accuracy of various LLMs on our MMedBench are shown, where our final model MMed-Llama 3 demonstrated the best performance among all existing open-source LLMs. d The comparison bar further details the gains in both multi-choice accuracy and ability of rationale generation, when comparing MMedLM 2 to InternLM 2, or comparing MMed-Llama 3 to Llama 3. Considering that the main difference between our models and their base models lies in the auto-regressive training on MMedC, such comparison highlights the importance of our contributed medical-specific multilingual language corpus.
@@
For auto-regressive training, we have developed a large-scale Multilingual Medical Corpus (MMedC), amassing over 25.5 billion medical-related tokens across six primary languages: English, Chinese, Japanese, French, Russian, and Spanish. This diverse dataset was compiled from four distinct sources: (i) we devised an automatic pipeline to filter medical-related content from the broad multilingual corpus, ensuring a focused and relevant dataset, (ii) we curated an extensive collection of medical textbooks in various languages, and converted them into texts with carefully designed pre-processing, e.g., Optical Character Recognition (OCR), heuristic data filtering, etc. We will share the name list of the books and the methodologies and tools for curation, (iii) to guarantee a wide-ranging encapsulation of medical knowledge, we incorporated texts from some open-source medical websites, enriching our corpus with authoritative and comprehensive medical information, (iv) we have integrated a number of existing small-scale medical corpus datasets, further enhancing the breadth and depth of ours. To our knowledge, MMedC represents the first endeavor to construct a corpus specifically focused on the multilingual medical domain.
@@
As for benchmark curation, we start by aggregating existing medical multiple-choice QA datasets across the six languages as of MMedC. We further augment them with rationale content by using GPT-4, enriching the datasets with explanations that support the correct answers. Consequently, our enriched dataset encompasses 53,566 QA pairs across six languages, uniquely offering both multi-choice QA and accompanying rationale reasoning. This extensive collection spans 21 medical fields, including but not limited to Internal Medicine, Biochemistry, Pharmacology, and Psychiatry, among others, termed as the Multilingual Medical Benchmark (MMedBench). We divide it into 45,048 training pairs and 8,518 testing pairs. The training split enables to finetune LLMs after domain-specific continues training. We utilize the entire test set, comprising 8,518 QA pairs, to evaluate the accuracy of multi-choice question answering. To further examine the models’ reasoning ability, we select a subset of 1,136 QA pairs, each accompanied by manually verified rationale sentences, serving as a more specialized benchmark for reasoning evaluation.
@@
At the evaluation phase, we conducted comprehensive benchmarking across eleven existing LLMs with multilingual support, including, GPT-3.5, GPT-4, Gemini-1.0 pro, BLOOM, InternLM, InternLM 2, MedAlpaca, ChatDoctor, PMC-LLaMA, Mistral, BioMistral, MEDITRON, Llama 2 and Llama 3, alongside the LLMs further trained with MMedC. These models were evaluated across three different settings: zero-shot, parameter-efficient fine-tuning (PEFT), and full fine-tuning. Given the complexity of evaluating rationale quality, which demands an assessment of long sentence semantic integrity, in addition to leveraging mainstream automated metrics, we also incorporate human rating scores in our analysis. This dual approach not only provides a comprehensive measure on each model’s performance, but also enables us to scrutinize the correlation between automated metrics and human judgment. Through this analysis, we identify the most reliable metric for extended comparisons, thereby enriching the methodology for evaluating reasoning ability in large language models.
@@
In our experiments, models that underwent further auto-regressive training on the MMedC consistently demonstrate enhanced performance, thereby underscoring the value and effectiveness of our compiled multilingual corpus. Our final model MMed-Llama 3 demonstrates the best performance on both multilingual and English-only benchmarks. We will publicly release our dataset (except for the license-restricted books for which we will provide a name list), codebase, and trained models to foster future research. In addition, we recognize the significance of robust evaluation metrics, particularly for the generation of medical texts that often involve complex, long sentences. To this end, we will also release detailed human rating results for individual cases.
@@ 2 Results
Here, we start by presenting the statistics of our constructed datasets. Then, we evaluate various LLMs on MMedBench on multi-choice questions and rationale ability, as well as verifying the effectiveness of MMedC. Lastly, we conduct a series of ablation studies to investigate the impact of each dataset component.
Data Statistics
@@
We first present the multilingual medical corpus (MMedC), which refers to a multilingual medical corpus containing over 25.5B tokens, acquired mainly from four sources, i.e., filtering medical-related content from the general large-scale multilingual corpus, medical textbooks, medical websites and existing small-scale corpus. The main statistic results are plotted in Figure 2.
@@
Figure 2: Statistic results on MMedC. a The Distribution of languages included in MMedC around the world. The map shows our collected corpora can cover most main countries worldwide. b The Token distribution for each language. The bar plot shows the detailed token number for different languages. c The Contributions of four sources to six languages for our MMedC. The Sankey diagram shows how the four considered data sources contribute for different languages, i.e., filtering content, medical textbooks, medical websites and small-scale corpus.
@@
Then, to better evaluate the performance of multilingual medical models, we further present a comprehensive multilingual medical Question and Answering Benchmark (MMedBench). We start by delving into its core attributes, including the total number of training and testing cases, the distribution of answer options, and the average length of question-and-answer tokens. Figure 3a illustrates these fundamental characteristics, highlighting that MMedBench often includes questions with multiple correct options, that brings complexity for models to navigate. Additionally, the answers contain rationale sections averaging 200 tokens each. This substantial token count serves two purposes: it helps to train language models by exposing them to extended reasoning passages, but also in evaluating their ability to generate and understand lengthy, complex reasoning statements.
@@
In our detailed exploration of MMedBench, we categorize each question into one of the 21 medical topic categories using GPT-4. These categories include Internal Medicine, Biochemistry, Pharmacology, Psychiatry, Microbiology, Physiology, Pathology, Immunology, Obstetrics and Gynecology, Public Health, Hematology, Surgery, Emergency Medicine, Orthopedics, Neurology, Anatomy, Medical Genetics, Radiology, Dermatology, and Endocrinology. This categorization has been rigorously verified by at least two clinicians, to ensure its comprehensiveness, and covering the breadth of medical disciplines.
@@
Figure 3b showcases the diversity of our multilingual benchmark, spanning a wide array of medical questions from foundational clinical medicine to specialized areas such as pharmacology and public health, with a pronounced emphasis on areas like Internal Medicine and Biochemistry. This underlines the effectiveness of the benchmark, to assess models’ ability on recognizing and processing a broad spectrum of medical inquiries effectively.
@@
Figure 3: Statistic results on MMedBench. a The bar plot shows the foundation statistic number on the train and test set of MMedBench. The term "Avg. tokens" represents the mean token length per sample across various compositions in it. “Rationale” denotes the rationale sentences in answer. “Option” denotes the option descriptions in choice list and “question” denotes the question sentences. Then the term “Prop. of multi-option” denotes the proportion of the question with multiple correct options and “Prop. of single-option” denotes the proportion of those with one options in answer. The final term “Number of QA pairs” denotes how many QA pairs are in train or test split. b The statistic histogram shows the topics distribution in the test split of MMedBench, covering a wide range of medical aspects, ranging from general and specialized medicine to basic medical sciences. This allows MedQA to comprehensively measure the performance of medical models.
Evaluation on MMedBench
@@
In this section, we present a comprehensive benchmark of the foremost LLMs using our MMedBench under zero-shot, PEFT, and full fine-tuning settings. Our evaluation focuses on two aspects of model performance: the accuracy in multiple-choice questions and the models’ ability to generate rationales. The evaluated LLMs can be categorized into four distinct classes, i.e., closed-source LLMs, popular open-source LLMs, medical-specific open-source LLMs, and those further undergone training on our MMedC. The latter three can all be categorized into open-source LLMs.
@@
Initially, our analysis focuses on the state-of-the-art, proprietary closed-source LLMs developed by OpenAI and Google, specifically GPT-3.5, GPT-4, and Gemini-1.0 pro. These models are examined through their publicly available online API solely in a zero-shot setting, as they are not accessible for any further training. However, note that, as the training data for these closed-source models are confidential, it is difficult to judge whether they are really “zero-shot”. Following this, our evaluation encompasses a range of open-source LLMs such as Mistral, InternLM 2 and Llama 3. We observe that the response from these open-source LLMs is relatively poor, making it difficult to draw effective comparisons (check Supplementary Material F for more zero-shot fail cases) in the zero-shot setting. We therefore compare them in fine-tuning settings (PEFT and full fine-tuning). Among these, we make a further distinction between general LLMs and those specifically tailored for the medical domain. Finally, we evaluate models that have undergone further training on our proposed corpus, named MMedLM (based on InternLM), MMedLM 2 (based on InternLM 2) and MMed-Llama 3 (based on Llama 3). These models are uniquely augmented with domain-specific knowledge by auto-regressive training on MMedC.
@@
We first evaluate models on Multilingual Multiple-choice Question & Answering tasks. As illustrated in Table 1, medical-specific Large Language Models (LLMs) generally exhibit high accuracy scores in English, yet their performance significantly declines in languages other than English. Notably, the finetuned PMC-LLaMA achieved an English accuracy score of 47.53, despite outperforming its contemporaneous counterparts, falls behind the GPT models significantly. Later, with the deployment of more advanced foundational models, open-source models started to bridge the gap with the GPT series, for instance, Mistral, InternLM 2, Llama 3, once fine-tuned on the train set of MMedBench, recorded average accuracy scores of 60.73, 58.59 and 62.79, respectively, surpassing all predecessors of comparable scale. Enhanced performance is also observed after additional auto-regressive training on our proprietary MMedC dataset. Specifically, our final model, MMed-Llama 3, demonstrated significant improvements over it’s counterparts without further training on MMedC, for example, 67.75 (MMed-Llama 3) vs. 62.79 (Llama 3) under full fine-tuning evaluation. A similar observation also holds for the PEFT setting, i.e., later LLMs performs better and training on MMedC brings a significant gain. As a result, MMed-Llama 3 refers to the most competitive open-source model with 8B parameters in proximity to GPT-4’s accuracy of 74.27.
[b]
@@
Parameter-efficient Fine-tuning (PEFT) Evaluation
@@
MMedLM 2 (Ours) 7B - trainset 61.7 80.01 61.81 52.09 80.47 67.65 67.30
MMed-Llama 3 (Ours) 8B - trainset 66.06 79.25 61.81 55.63 75.39 68.38 67.75
@@
We offer detailed instructions that prompt the model to outline its analytical process for delivering the final answer, enabling a clear assessment of its reasoning capabilities. The performance is then meticulously evaluated using a variety of metrics. Specifically, the ROUGE-1 and BLEU-1 scores are presented in Table 2. Additionally, results for other metrics are detailed in Supplementary Material E providing a comprehensive view of the model’s performance across diverse evaluation frameworks.
[b]
@@
BioMistral 45.85/41.34 43.12/44.75 38.76/44.46 41.82/39.41 27.73/18.80 45.52/41.07 40.46/38.31
Llama 3 46.33/ 41.73 47.09/ 47.44 46.24/ 50.43 43.13/ 40.69 30.89/ 22.22 47.14/ 42.70 43.47/ 40.87
MMedLM (Ours) 41.63/ 38.83 44.30/ 46.38 38.61/ 46.90 37.54/ 37.78 19.99/ 21.28 40.79/ 38.77 37.14/ 38.32
MMedLM 2 (Ours) 47.07/ 41.51 47.15/ 48.36 47.90/ 52.24 43.22/ 41.36 27.81/ 25.70 46.17/ 42.64 43.22/ 41.97
MMed-Llama 3 (Ours) 46.56/ 41.57 47.12/ 47.71 48.10/ 53.18 43.62/ 40.97 33.92/ 24.87 47.67/ 43.32 44.50/ 41.94
@@
BioMistral 47.96/ 42.16 49.76/ 49.33 49.73/ 52.12 46.34/ 41.64 34.20/ 24.27 47.57/ 41.11 45.93/ 41.77
Llama 3 48.74/ 43.66 49.44/ 49.426 51.97/ 53.98 47.11/ 42.49 34.73/ 25.07 48.59/ 42.44 46.76/ 42.84
MMedLM (Ours) 47.37/ 41.98 48.68/ 49.28 48.95/ 52.34 45.39/ 41.41 33.24/ 24.67 46.68/ 41.35 45.05/ 41.84
MMedLM 2 (Ours) 50.02/ 44.77 51.39/ 51.78 54.79/ 57.10 49.04/ 45.30 37.49/ 28.18 50.14/ 44.59 48.81/ 45.29
MMed-Llama 3 (Ours) 47.61/ 42.47 49.96/ 49.36 52.89/ 55.06 47.92/ 42.85 36.31/ 26.67 48.61/ 43.41 47.21/ 43.29
@@
Specifically, from the test set of MMedBench, we randomly selected 50 test cases per language, alongside outcomes generated by six notable models: MMed-Llama 3 (ours), Llama 3, InternLM 2, BioMistral, MEDITRON, GPT-3.5. The sequence of samples and corresponding model outputs are randomized to prevent bias. The evaluation panel, comprising five post-graduate students from the medical school of Shanghai Jiao Tong University and Peking Union Medical College, was instructed to rank the outputs based on accuracy, reasoning ability, and internal knowledge. To facilitate accurate assessment, manually verified references were also provided. Rankings were quantitatively assigned, with the highest rank awarded a score of 6 and the lowest a score of 1, thereby quantifying the quality of each model’s output. In parallel, we leveraged GPT-4 as an additional evaluator, assigning it the role of a judge to rank the outputs. Further details on GPT-4’s evaluation method are available in Supplementary Material A.
@@
Figure 4a illustrates the comparative analysis of model performances through relative ratings. Notably, MMed-Llama 3 achieved the highest scores in both human (4.10) and GPT-4 (4.73) evaluations, aligning with its superior performance as indicated by the automatic machine metrics. It is particularly worth highlighting that MMed-Llama 3 can significantly outperform the other models in the GPT-4 rating, surpassing the second best model InternLM 2 by 0.89 rating scores. Interestingly, GPT-3.5 received a lower human rating of 2.37, suggesting that the evaluators’ preferences might be influenced by the brevity of the responses. Comprehensive rating results for each language and model are detailed in Supplementary Material E.
@@
In addition to comparing different LLMs, our study delves into the correlation between various automatic evaluation metrics and human preferences. This correlation analysis enables us to identify the most effective automatic metric for benchmarking purposes, thereby potentially eliminating the need for resource-intensive human evaluations in future research. We employ the Kendall rank correlation coefficient to measure the agreement between the rankings of each model’s generated rationales by automatic metrics and human evaluations. The findings, illustrated in Figure 4b, indicate that GPT-4’s evaluation results have the highest correlation with human judgments, with a $\tau$ value of 0.660. However, it’s important to note that GPT-4’s ratings, while highly correlated, are relative and not easy to scale for evaluating newly introduced models. Among the absolute automatic metrics, BERT Score emerged as the most reliable indicator, demonstrating a $\tau$ value of 0.538. Consequently, we advocate for the use of Bert Score as benchmark for assessing the rationale capabilities of newly introduced LLMs on MMedBench in subsequent studies.